OpenStudy (anonymous):

11. In how many ways can the 11 identical horses on a carousel be painted so that three are brown, three are white, and five are black?

OpenStudy (anonymous):

doing the brown, then white, then black 3 Brown starting with 11 possibilities 11*10*9 / (1*2*3) = 165 3 White with 8 remaining possibilities 8*7*6 / (1*2*3) = 56 5 black with the 5 remaining possibilities = 1 5*4*3*2*1 / (1*2*3*4*5) 165 * 56 * 1 = 9240 Doing the black first, we should come up with the same answer Balck, brown then white 11*10*9*8*7 / (1*2*3*4*5) = 462 6*5*4 / (1*2*3) = 20 3*2*1 / (1*2*3) = 1 462 * 20 *1 = 9240 9240